OpenStudy (aric200):

Which is the equation for the line that passes through (−4, −3) and (−2, 3)? y = 3x + 15 y = 3x + 9 y = -3x - 15 y = -3x - 9

OpenStudy (aric200):

Looking for help to or answer with explanation

OpenStudy (anonymous):

for problems like this.. I would just copy and paste them into the input line on geogebra to cross check them..

OpenStudy (anonymous):

Find the gradient of the line using the formula, m= y2-y1/x2-x1 than to find the equation use this formula, y-y1= m(x-x1)